# LUT-LLM: Efficient Large Language Model Inference with Memory-based Computations on FPGAs

### v1

@@
Abstract.
The rapid development of large language models (LLM) has greatly enhanced everyday applications. While cloud-based LLM services often dominate attention due to their massive processing capabilities, efficient single-batch inference remains a critical research direction for enabling on-device intelligence. In this realm, FPGA-based accelerators, with their micro-architectural flexibility for fine-grained data control, are well recognized for achieving superior speed and energy efficiency compared to GPUs. However, recent GPU-specific optimizations have diminished this advantage. When limited to arithmetic-based computation, FPGAs can underperform GPUs due to their comparatively fewer computational resources.
@@ Abstract.
To address this challenge, we exploit a key advantage of FPGAs over GPUs: abundant distributed on-chip memory embedded among computational units. We believe that shifting language model inference from arithmetic-based to memory-based computations through table lookups can improve the efficiency on FPGAs to compete with GPUs. However, existing memory-based accelerators are inefficient or unable to scale to large models, and fail to capture the execution properties of language models. This paper introduces LUT-LLM, the first FPGA accelerator for 1B+ language model inference with memory-based computation, leveraging vector quantization. We construct a performance model, evaluate multiple quantization schemes, and identify activation-weight vector co-quantization as the most effective approach. To support this scheme for LLM inference, LUT-LLM features (1) bandwidth-aware parallel centroid search to reduce decoding latency, (2) efficient 2D table lookups for activation-weight co-quantization, and (3) a spatial-temporal hybrid design to reduce data caching for a higher throughput table lookup. We prototype LUT-LLM for a customized Qwen 3 1.7B model on the AMD V80 FPGA, achieving a $1.66\times$ lower end-to-end latency than the AMD MI210 GPU and a $1.72\times$ higher energy efficiency than the NVIDIA A100 GPU. LUT-LLM can scale to a 32B model with a $2.16\times$ better energy efficiency than the A100.
Vector Quantization, Language Model Inference, Table Lookup
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Hardware Hardware acceleratorsccs: Computer systems organization Neural networks
1. Introduction
The advancement of LLMs has enabled a wide range of applications such as conversational chatbots (dam2024complete; oruche2025survey), coding agents (zhang2024codeagent; li2024codetree), and deep research (huang2025deep) that assists our daily lives. Mainstream LLM services are deployed through cloud-based infrastructures with clusters of data center GPUs and serve multiple user requests concurrently. While effective at scale, this paradigm faces limitations when tackling edge scenarios such as personal AI assistants (cheng2024autopal; lee2024towards), smart home devices (qualcomm2025smarthomes), and interactive robotics (ayub2024interactive), where samples cannot be batched to fully utilize the resources. Consequently, efficient single-batch inference has emerged as a crucial research direction to enable on-device intelligence.
@@ 1. Introduction
FPGAs have been widely recognized as promising platforms for single-batch language model inference, offering competitive speed and energy efficiency relative to GPUs (chen2024understanding; chen2024allo; zeng2024flightllm). Nevertheless, recent advancements in GPU architectures and software optimization have narrowed this performance gap. In particular, serving a Llama 2 7B model on a single NVIDIA A100 GPU obtains a $2.38\times$ throughput improvement and a $3.27\times$ energy efficiency gain when FlashAttention (dao2023flashattention), FlashDecoding (hong2023flashdecoding++), and quantization with GPTQ (frantar2022gptq) are enabled. Under these settings, GPU inference is $2.00\times\sim 6.37\times$ faster with similar or better energy efficiency than state-of-the-art (SoTA) FPGA accelerators (chen2024allo; zeng2024flightllm). These results highlight the growing competitiveness of GPUs in efficient inference and suggest that sustaining FPGA’s advantages will require algorithmic and architectural innovations.
@@
Figure 1. Motivations and challenges of memory-based computation for LLM inference on FPGA, with the corresponding solutions as the technical contributions in LUT-LLM.
@@
When competing with GPUs, the major challenge is that FPGAs often have fewer computational resources than GPUs’ process node (amd2023alveou280; nvidia2020v100; amd2024alveov80; nvidia2021a100). This leads to slow input processing and underutilized memory bandwidth because FPGAs spend more time on computation than on memory access. However, FPGAs typically provide larger amounts of distributed on-chip memory units compared to GPUs. Depicted in Figure 1, the AMD V80 (amd2024alveov80) integrates $14.9\times$ more on-chip memory units with a $2.5\times$ larger capacity than the NVIDIA A100 (nvidia2021a100). Therefore, shifting the computational paradigm from arithmetic-based operations to memory-based computations is promising. Prior memory-based accelerators (li2025lut; gerlinghoff2024table; andronic2025neuralut; miyasaka2024synthesis) show the efficiency on neural networks within 10M parameters, but face two challenges when deploying language models (Figure 1):
@@
  • Scaling inefficiency and infeasibility (C1): direct deployment of existing methods can lead to inefficiency due to higher memory and computation cost compared to conventional arithmetic approaches, or infeasible with limited on-chip resources.
@@
  • Additional Complexity of LLMs (C2): prior memory-based neural accelerators overlook the additional complexity of LLM computations, where centroid search is hard to pipeline in decoding, lookup table access is limited by on-chip memory ports, and interactions between linear layers and other components further hinder data movement efficiency.
@@
In this paper, we propose LUT-LLM  [^1] In this paper, “LUT” in “LUT-LLM” and the architecture diagram refers to a general lookup table, whereas elsewhere it denotes the LUT in FPGA fabric., the first FPGA accelerator for 1B+ language models with memory-based computations. LUT-LLM replaces the computations in linear layers with table lookups that store pre-computed dot product results. To enable efficient table construction, we apply vector quantization (gray1984vector), where multiple values in matrices are quantized into low-bit indices. To resolve C1, we first develop a performance model for model inference with memory-based computation and vector quantization to understand the scaling (Section 3). We innovatively show that vector quantizing both weights and activations with pre-computed 2D lookup tables is feasible and achieves the best performance. From this foundation, we design the LUT-LLM accelerator with the following three features that resolves C2 (Section 4):
@@
  • Bandwidth-aware parallel centroid search: LUT-LLM finds the closest centroid of every input vector to access the corresponding entry in the lookup tables. Unlike a naive single pipeline (li2025lut) or a complete binary reduction tree architecture, LUT-LLM employs a hybrid design with multiple parallel pipelines and a smaller reduction tree. The parallelization size and pipeline depth are co-designed with the memory bandwidth to ensure the maximum throughput of the compute engine by fully hiding the centroid search latency.
@@
  • Spatial-temporal hybrid design: LUT-LLM adopts a hybrid execution strategy, following a design principle similar to InTAR (he2025intar) and SSR (zhuang2024ssr). Specifically, the attention layer is implemented in a dataflow manner, whereas linear layers are executed sequentially. This design spares $14\%$ on-chip buffers from attention to the linear layer for more parallel accesses, increasing the throughput.
We implement a prototype of LUT-LLM targeting a customized Qwen-3 1.7B model on an AMD V80 FPGA. Comparing with GPUs evaluated using vLLM (kwon2023efficient) with the model quantized to INT8 via GPTQ, experimental results demonstrate that our design is $1.66\times$ faster and is $4.1\times$ more energy efficient compared to AMD MI210 (amd2025mi210). It is also $1.72\times$ more energy efficient than the NVIDIA A100 GPU. Extending LUT-LLM for the Qwen 3 32B model can have $2.16\times$ energy efficiency boost than the A100.
2. Background
2.1. Language Model Acceleration
2.1.1. Model Architecture
Most modern language models (dubey2024llama; touvron2023llama) adopt the transformer architecture (vaswani2017attention) with stacked blocks of attention and feedforward networks (FFNs). While the architecture remains consistent, recent designs (yang2025qwen3; dubey2024llama) replace multi-head attention (MHA) with grouped-query attention (GQA) (ainslie2023gqa).
Given input $X\in\mathbb{R}^{L\times D}$ with sequence length $L$ and hidden size $D$, GQA projects $X$ into queries, keys, and values. Unlike MHA, GQA maps multiple queries to each key-value pair, reducing cache size while maintaining accuracy. For $h$ query heads and $g$ heads per group, head $i$ is computed as:
(1) $\begin{split}Q_{i}=XW_{Q_{i}},\;K_{i/g}=XW_{K_{i/g}},\;V_{i/g}=XW_{V_{i/g}},\\ O_{i}=\text{softmax}(Q_{i}K_{i/g}^{T}/\sqrt{d})V_{i/g},\end{split}$
where $W_{Q_{i}},W_{K_{i/g}},W_{V_{i/g}}\in\mathbb{R}^{D\times d}$ and $d=D/h$, with keys and values shared within groups. Concatenating $\{O_{i}\}_{i=1}^{h}$ into $O\in\mathbb{R}^{L\times D}$ and applying a linear projection produces the FFN input:
(2) $X^{\prime}=(f(OW_{1})\odot OW_{2})W_{3},$
where $f(\cdot)$ is a nonlinearity, $\odot$ denotes element-wise multiplication, and $W_{1},W_{2},W_{3}$ are FFN parameters.
Language model inference proceeds in two stages (chen2024understanding): the prefill stage encodes the full input sequence to generate the first token, while the decode stage autoregressively produces subsequent tokens conditioned on prior outputs.
2.1.2. Prior LLM Accelerators
@@
Previous works (chen2024allo; zeng2024flightllm; he2025intar; hong2022dfx) develop customized accelerators for model inference with arithmetic-based operations. Some accelerators, such as FlightLLM and DFX (hong2022dfx; zeng2024flightllm), sequentially execute operations with various sparsification and quantization techniques. Allo (chen2024allo) builds a Pythonic framework to easily construct dataflow accelerator for LLMs, and InTAR (he2025intar) proposes a spatial-temporal hybrid design for efficient on-chip management of LLMs.
2.2. Weight Vector Quantization
Figure 2. Linear projection with weight vector quantization.
@@ 2.2. Weight Vector Quantization
Conventional quantization compresses scalars into low-bit representations via scaling and shifting (nagel2020up; lin2024awq; xiao2023smoothquant; frantar2022gptq). In contrast, vector quantization (VQ) (gray1984vector) encodes groups of $v$ elements as indices to representative centroids. As seen in Figure 2, in the preprocessing stage, the quantizer ➀ partitions the matrix into length-$v$ vectors, clusters them into $c_{w}$ centroids forming a codebook, and ➁ replaces each vector with the index of its closest centroid. The resulting weight centroid index table stores indices at a much lower bitwidth. For example, with $v=2$, vector $(1,3)$ is encoded as index 0, mapped to centroid $(2,3)$.
@@
For language models, VQ is applied to weight matrices to reduce memory cost (van2024gptvq; tseng2024quip; tseng2024qtip; liu2024vptq). An $M\times D$ matrix is divided into groups of $G$ vectors, each with its own codebook and index table (e.g., $G=4$ in Figure 2). At inference, the device ➂ loads codebooks and indices, reconstructs the weights by centroid lookup, and ➃ multiplies the recovered matrix with activations. Compared to scalar quantization, VQ provides non-uniform compression (van2024gptvq), yielding higher flexibility and representational power, and thus superior accuracy at the same bitwidth. However, applying VQ to weights requires significant floating-point operations that are costly on FPGAs.
2.3. Memory-based Computation
2.3.1. Benefits of Memory-based Computation
@@
Memory-based computation replaces arithmetic by loading pre-computed outputs from memory, making it well-suited for on-chip memory-rich platforms such as FPGAs. It reduces resource usage and latency since a multiply-accumulate (MAC) operation — requiring two reads, one multiplication, and one addition in arithmetic-based computations — can be replaced by a single memory read. This also lowers energy: a memory-based MAC consumes only 3.8 pJ in 7 nm, $2.4\times$ less than its arithmetic counterpart (jouppi2021ten). Moreover, memory-based LLMs avoid online dequantization because values are pre-computed in full precision. The pre-compute value can be further quantized per-tensor with low loss (li2025lut), which is used in our design.
@@
Figure 3. Linear projection with activation vector quantization. Precomputed dot products between weight matrix and centroids in the codebook are stored in the lookup tables.
2.3.2. ML Accelerators with Memory-based Computations
@@
For LLMs, weight quantization alone is insufficient, as lookup tables for dynamic activations are prohibitively large. Recent works (li2025lut; tang2023lut; wei2025t) therefore apply VQ to activations in linear layers. As illustrated in Figure 3, the quantizer collects activation samples during training. To preserve accuracy, methods such as LUT-NN (tang2023lut) ➀ learn codebooks with $c_{a}$ centroids, each through differentiable training and ➁ pre-compute dot products between weight vectors and centroids, storing them as lookup tables. For example, a weight vector $(1,3)$ multiplied by centroids $(6,2)$ and $(4,5)$ yields entries $12$ and $19$, producing $M\times D/v$ tables from an $M\times D$ weight matrix. At inference, the quantizer ➂ identifies the nearest centroid for each input vector, ➃ retrieves the corresponding table entry, and accumulates the results. For instance, input $(6,3)$ maps to centroid $(6,2)$ at index 0, retrieving $12$. LUT-DLA (li2025lut) is a recent work which targets an ASIC design. This work improves the learning and accelerator design for the framework activation VQ, which, while effective in CNNs, requires storing and loading large tables for LLMs, and, thus, limits the decode and short-context prefill performance (Section 3). In addition, its single-pipeline centroid search creates bottlenecks in the decode stage (Section 4).
Several other works, including TLMAC (gerlinghoff2024table) and NeuraLUT (andronic2025neuralut), consider using memory-based computations to replace MAC operations or simple neurons with FPGA soft logic. However, the limited resource makes them infeasible to scale to language models that compete with GPUs.
3. LUT-LLM Algorithm and Performance Modeling
In this section, we illustrate the performance model for language model inference with memory-based computations to guide our hardware design in Section 4.
3.1. Performance of Vector Quantized Linear Layer
@@
We start with modeling the performance of the linear layer, which is affected by the vector quantization, and then extend this to the complete language model. Table 1 illustrates the symbol utilized in the performance model for the FPGA resource assigned to the given linear layer and the quantization configurations. Here, we mainly show the key results for different quantization strategies, and leave the detailed derivation in Section 7.
Table 1. Symbols used in performance modeling
@@
$Op_{fp32}$ FP32 MACs per cycle per compute unit $c_{a}$ Centroids per activation codebook
$Op_{int8}$ INT8 MACs per cycle per compute unit $M\times D$ Weight matrix dimension
$C$ Off-chip bandwidth (bytes/cycle) $L$ Sequence length
@@
(4) $\begin{split}T_{lat}&=\frac{MD(\log(c_{w})/v+32/Gv)}{N_{p}b_{p}}+\frac{MDL}{\min(N_{c}Op_{fp32},N_{p}b_{p}/32)}\end{split}$
When overlapping off-chip memory access with on-chip computations by double buffering, the overall latency is $\max(T_{mem},T_{lat})$. For example, with $M=512$, $D=32$, $G=256$, $c_{w}=16$, $v=2$, $L=1$, $C=64$, and when we assign 16 ports of 32-bit wide memory and 256 FP32 compute units, $T_{mem}=66$ and $T_{lat}=1090$, resulting in $1090$ cycles for overall latency.
@@
(6) $\begin{split}T_{tl}&=SML/\min(SM,N_{p}b_{p}/8)\\ &+\frac{SML}{\min(SM,(N_{c}-Sc_{a}v/Op_{fp32})Op_{int8},N_{p}b_{p}/8)}\end{split}$
@@
(7) $T_{lat}=\min_{S}(\frac{D}{S}\max(\log(c_{a})+L-1,T_{tl}))$
With the same setting as the previous example and $c_{a}=64$, $T_{mem}=8256$ and $T_{lat}=512$, leading to $8256$ cycles for overall latency.
Figure 4. Linear projection with both activation and weight vector quantization with 2D lookup tables.
Prior works (van2024gptvq; tang2023lut) apply either weight-only or activation-only vector quantization. We propose that these approaches can be extended to activation-weight co-quantization. As illustrated in Figure 4, the process involves ➀ generating codebooks for weights and activations, ➁ pre-computing dot products between them to form 2D lookup tables, ➂ finding the closest centroid for each activation vector, and ➃ accessing the entry based on the activation and weight centroid indices for accumulation. In the figure, vector $(6,3)$ is closer to $(6,2)$ at index 0. As the first weight centroid index is 0, we access the lookup table entry $(0,0)$ with value $18$.
@@
(9) $\begin{split}T_{tl}&=SML/G\min(SM/G,N_{p}b_{p}/8)\\ &+\frac{SML}{\min(SM,(N_{c}-Sc_{a}v/Op_{fp32})Op_{int8},N_{p}b_{p}/8)}\end{split}$
(10) $T_{lat}=\min_{S}(\frac{D}{S}\max(\log(c_{a})+L-1,T_{tl}))$
@@
3.2. Roofline Analysis and Comparison
We extend the performance model of vector-quantized linear layers to a complete transformer architecture following (chen2024understanding), with additional modeling for newly introduced computation blocks in modern language models (e.g., GQA, SwiGLU, etc.). Following prior works (wei2025t; wang2023bitnet; wang2024bitnet; lin2024qserve), attentions and non-linear operations require floating point precision for accuracy. As an example, Figure 5 depicts the roofline analysis of deploying the Qwen 3 1.7B model on the AMD V80 FPGA with a naive memory-based computation employing activation vector quantization like (li2025lut; wei2025t) and arithmetic-based computation with FP16 quantization. The quantization configuration is: $c_{a}=64$, $v=2$, INT8 lookup tables, and FP32 codebooks.
Figure 5. Roofline model of Qwen 3 1.7B on AMD V80 FPGA with activation VQ for memory-based compute and arithmetic-based compute in FP16. Memory-based compute can attain higher throughputs for long sequence, but is highly memory bounded for decoding.

### v2

# LUT-LLM: Efficient Language Model Inference with Memory-based Computations on FPGAs

@@
Abstract
@@
The rapid development of large language models (LLM) has greatly enhanced everyday applications. While many FPGA-based accelerators, with flexibility for fine-grained data control, exhibit superior speed and energy efficiency compared to GPUs, recent GPU-specific optimizations have diminished this advantage. When limited to arithmetic-based computation, FPGAs often underperform GPUs due to their comparatively fewer computational resources. To address this challenge, we exploit a key advantage of FPGAs over GPUs: abundant distributed on-chip memory embedded among computational units. We believe that shifting LLM inference from arithmetic-based to memory-based computations through table lookups can improve the efficiency on FPGAs to compete with GPUs. However, existing methods are inefficient or unable to scale and deploy language models due to algorithm and architecture design limitations. This paper introduces LUT-LLM [^1] Github: https://github.com/LUT-FPGA/LUT-LLM, the first FPGA accelerator that deploy 1B+ language model with memory-based computation, leveraging vector quantization. We construct a performance model, evaluate multiple quantization schemes, and identify activation-weight vector co-quantization as the most effective approach. To support this scheme, LUT-LLM features (1) bandwidth-aware parallel centroid search to reduce decoding latency, (2) efficient 2D table lookups, and (3) a spatial-temporal hybrid design to reduce data caching for a higher throughput table lookup. We develop a training recipe that converts existing models to support table lookups with high accuracy and prototype LUT-LLM for Qwen 3 1.7B model on the AMD V80 FPGA, reducing arithmetic operations by $4\times$ and achieving a $1.10\sim 3.29\times$ faster generation speed and a $3.05\sim 6.60\times$ higher energy efficiency than GPUs.
I Introduction
@@
The advancement of LLMs has enabled a wide range of applications such as conversational chatbots [13, 42], coding agents [64, 31], and deep research [25] that assist our daily lives. Recently, FPGA have been widely recognized as promising platforms for efficient single-batch LLM inference compared with GPUs [9, 10, 63]. This covers edge scenarios such as personal AI assistants [11, 29], smart home devices [44], and interactive robotics [7]. Nevertheless, the advancements in GPU kernel and algorithm optimization have narrowed this performance gap. In particular, serving a Llama 2 7B model on a single NVIDIA A100 GPU obtains a $2.38\times$ throughput improvement and a $3.27\times$ energy efficiency gain when FlashAttention [14], FlashDecoding [23], and quantization with GPTQ [15] are enabled. Under these settings, GPU inference is $2\times\sim 6.37\times$ faster with similar or better energy efficiency than state-of-the-art FPGA accelerators [10, 63]. These results highlight the growing competitiveness of GPUs in efficient inference and suggest that sustaining FPGA’s advantages will require algorithmic and architectural innovations.
@@
Figure 1: Motivations and challenges of memory-based computation for LLM inference on FPGA, with the corresponding solutions as the technical contributions in LUT-LLM.
@@
When competing with GPUs, the major challenge is that FPGAs often have fewer computational resources than GPU [3, 39, 5, 40]. This leads to slow input processing and underutilized memory bandwidth when FPGAs spend more time on computation than on memory access. However, FPGAs typically provide larger amounts of distributed on-chip memory units: the AMD V80 (7nm) [5] integrates $14.9\times$ more on-chip memory units with a $2.5\times$ larger capacity than the NVIDIA A100 (7nm) [40]. Therefore, shifting the computational paradigm to memory-based computations is promising. Prior memory-based accelerators [30, 17, 6, 37] show the efficiency on neural networks within 10M parameters, but face two challenges when deploying language models (Figure 1):
@@
  • Scaling inefficiency and infeasibility (C1): direct deployment of existing methods that convert models for table lookups [30, 35, 50] can lead to inefficiency due to higher memory and computation cost compared to conventional arithmetic approaches, or be infeasible with limited on-chip resources.
@@
  • Additional Complexity of LLMs (C2): prior memory-based neural accelerators overlook the additional complexity of LLM accelerations, where centroid search is hard to pipeline in decoding, lookup table access is limited by on-chip memory ports, and interactions between linear layers and other components further hinder data movement efficiency.
@@
In this paper, we propose LUT-LLM, the first FPGA accelerator targeting 1B+ parameter LLMs using memory-based computation. LUT-LLM replaces conventional linear layers with table lookups over pre-computed dot-product results. To enable efficient table construction, we adopt vector quantization [18], which maps multiple values to low-bit indices. This design reduces the number of operations and shortens per-operation latency, enabling competitive performance compared to GPUs that often underutilize available memory bandwidth [63]. To address C1, we develop a performance model for LLM inference with memory-based computation and vector quantization to understand the scaling (Section III). We innovatively show that vector quantizing both weights and activations achieves the best performance while maintaining comparable accuracy. From this foundation, we design the LUT-LLM accelerator with three features that resolve C2 (Section IV):
@@
  • Bandwidth-aware parallel centroid search: LUT-LLM finds the closest centroid of every input vector to access the corresponding entry in the lookup tables. Unlike a naive single pipeline [30] or a complete binary reduction tree architecture, LUT-LLM employs a hybrid architecture with multiple parallel pipelines and a smaller reduction tree. The parallelization size and pipeline depth are co-designed with the memory bandwidth to maximize throughput while hiding search latency.
@@
  • Spatial-temporal hybrid design: LUT-LLM adopts a hybrid execution strategy, following a design principle similar to InTAR [22] and SSR [65]. Specifically, the attention layer is implemented in a dataflow manner, whereas linear layers are executed sequentially. This design spares $14\%$ on-chip buffers from attention to the linear layer for more parallel accesses.
We develop a training recipe for vector-quantized model for our scheme to convert and deploy Qwen 3 1.7B on LUT-LLM prototyped with AMD V80 FPGA. Compared with GPUs at the same technology node, our design achieves a $1.10-3.29\times$ speedup and $3.05-6.60\times$ higher energy efficiency than AMD MI210 and NVIDIA A100.
II Background
II-A Language Model Acceleration
@@
Most LLM accelerators adopt the transformer architecture [51], which contains a series of attention blocks implemented in grouped-query attention (GQA) [2] and feedforward networks (FFN) with residual connections [9]. Previous works [10, 63, 22, 24] develop customized transformer accelerators with arithmetic-based operations. FlightLLM and DFX [24, 63] sequentially execute operations with various sparsification and quantization techniques. Allo [10] and StreamTensor [61] are dataflow accelerator design frameworks that can be used for LLMs, and InTAR [22] proposes a spatial-temporal hybrid design for efficient on-chip management of LLMs.
II-B Weight Vector Quantization
Figure 2: Linear projection with weight vector quantization.
@@ II-B Weight Vector Quantization
Conventional quantization compresses scalars into low-bit representations via scaling and shifting [38, 32, 59, 15]. In contrast, vector quantization (VQ) [18, 1] encodes groups of $v$ elements as indices to representative centroids. As seen in Figure 2, in the preprocessing stage, the quantizer ➀ partitions the matrix into length-$v$ vectors, clusters them into $c_{w}$ centroids forming a codebook, and ➁ replaces each vector with the index of its closest centroid. The resulting weight centroid index table stores indices at a much lower bitwidth. For example, with $v=2$, vector $(1,3)$ is encoded as index 0, mapped to centroid $(2,3)$.
@@
For language models, VQ is applied to weight matrices to reduce memory cost [50, 48, 49, 35]. An $M\times D$ matrix is divided into groups of $G$ vectors, each with its own codebook and index table (e.g., $G=4$ in Figure 2). In inference, the device ➂ loads codebooks and indices, reconstructs the weights by centroid lookup, and ➃ multiplies the recovered matrix with activations. Compared to scalar quantization, VQ provides non-uniform compression [50], yielding higher representational power for superior accuracy at the same bitwidth.
II-C Memory-based Computation
@@
Memory-based computation replaces arithmetic by loading pre-computed outputs from memory, making it well-suited for on-chip memory-rich platforms such as FPGAs. It reduces resource usage and latency since a multiply-accumulate (MAC) operation (two reads, one multiply and one addition) can be replaced by a single memory read. This also lowers energy: a memory-based MAC consumes only 3.8 pJ in 7 nm, $2.4\times$ less than its arithmetic counterpart [27]. Moreover, memory-based LLMs avoid online dequantization because values are pre-computed in full precision. The pre-computed lookup table can be further quantized with low loss [30], used in our design.
@@
Figure 3: Linear projection with activation VQ. Precomputed dot products between weight matrix and centroids are stored in the lookup tables.
@@
Previous works, including TLMAC [17] and NeuraLUT [6], consider using memory-based computations for simple neurons with FPGA soft logic. However, the limited resource makes it infeasible to scale to language models that compete with GPUs. For LLMs, weight quantization in these methods requires large storage for offline lookup tables due to the wide range of activations. Recent works [30, 47, 57, 45] therefore apply VQ to activations in linear layers. As illustrated in Figure 3, the quantizer collects activation samples during training. To preserve accuracy, methods such as LUT-NN [47] ➀ learn codebooks with $c_{a}$ centroids through differentiable training and ➁ pre-compute dot products between weight vectors and centroids, storing them as lookup tables. For example, a weight vector $(1,3)$ multiplied by centroids $(6,2)$ and $(4,5)$ yields entries $12$ and $19$. During inference, the quantizer ➂ identifies the nearest centroid for each input vector, ➃ retrieves the corresponding table entry, and accumulates the results. For instances, input $(6,3)$ maps to centroid $(6,2)$ at index 0, retrieving $12$. LUT-DLA [30] is a hardware-algorithm co-design that improves model learning and accelerator design. Though effective in CNNs, it still requires non-trivial large storage for LLMs, which limits decode and short-context prefill performance. In addition, its single-pipeline centroid search creates bottlenecks in the decode stage.
III Algorithm and Performance Modeling
In this section, we illustrate the performance model of our new quantization scheme to guide our hardware design in Section IV. Detailed derivation is omitted to save space.
III-A Performance of Vector Quantized Linear Layer
@@
We start with modeling the performance of the linear layer, which is affected by the vector quantization, and then extend this to the complete language model. Table I illustrates the symbols utilized in the performance model. The equivalent quantization scheme is $\log(c_{w})/v$-bit for weight and $\log(c_{a})/v$-bit for activation. Here, we mainly show the key results for different quantization strategies, and leave the detailed derivation in Appendix VII.
TABLE I: Symbols used in performance modeling of linear layer
@@
$Op_{fp32}$ FP32 MACs/cycle/compute unit $c_{a}$ Centroids per activation codebook
$Op_{int8}$ INT8 MACs/cycle/compute unit $D,M$ Input/Output weight dimensions
$C$ Off-chip bandwidth (bytes/cycle) $L$ Activation (Sequence) length
@@
$\begin{split}T_{lat}&=\frac{MD(\frac{\log(c_{w})}{v}+\frac{32}{Gv})}{N_{p}b_{p}}+\frac{MDL}{\min(N_{c}Op_{fp32},\frac{N_{p}b_{p}}{32})}\end{split}$ (2)
When overlapping off-chip memory access with on-chip computations by double buffering, the overall latency is $\max(T_{mem},T_{lat})$. For example, with $M=512$, $D=32$, $G=256$, $c_{w}=16$, $v=2$, $L=1$, $C=64$, and when we assign 16 memory ports of 32-bit wide memory and 256 FP32 compute units, $T_{mem}=66$ and $T_{lat}=1090$, resulting in $1090$ cycles for overall latency.
@@
$\begin{split}T_{tl}&=SML/\min(SM,N_{p}b_{p}/8)\\ &+\frac{SML}{\min(SM,(N_{c}-Sc_{a}v/Op_{fp32})Op_{int8},N_{p}b_{p}/8)}\end{split}$ (4)
@@
$T_{lat}=\min_{S}(D\max(\log(c_{a})+L-1,T_{tl})/S)$ (5)
With the same setting as the previous example and $c_{a}=64$, $T_{mem}=8256$ and $T_{lat}=512$, leading to $8256$ cycles for overall latency. Notice that some table-lookup-based FPGA accelerators [17], while quantizing weights, can be modeled in this formula by replacing $c_{a}/v$ with $2^{b}$ and divide $D$ by quantization group size, where $b$ is the activation bitwidth.
Figure 4: Linear projection with both activation and weight vector quantization with 2D lookup tables.
From the analysis in this example, we observe that either weight-only or activation-only VQ applied in prior works [50, 47] incur high computation latency or require loading large tables. To resolve these issues, we extend these schemes to activation-weight co-quantization, which quantizes in coarser-grain but still maintain comparable accuracy as shown in Table III. As illustrated in Figure 4, the process involves ➀ generating codebooks for weights and activations, ➁ pre-computing dot products between them to form 2D lookup tables, ➂ finding the closest centroid for each activation vector, and ➃ accessing the entry based on the activation and weight centroid indices for accumulation. In the figure, vector $(6,3)$ is closer to $(6,2)$ at index 0. As the first weight centroid index is 0, we access the lookup table entry $(0,0)$ with value $18$.
@@
$\begin{split}T_{tl}&=SML/G\min(SM/G,N_{p}b_{p}/8)\\ &+\frac{SML}{\min(SM,(N_{c}-Sc_{a}v/Op_{fp32})Op_{int8},N_{p}b_{p}/8)}\end{split}$ (7)
$T_{lat}=\min_{S}(\tfrac{D}{S}\max(\log(c_{a})+L-1,T_{tl}))$ (8)
@@
III-B Analysis and Comparison